# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ Abstract.
We consider $m\times n$ rectangular matrices formed from Sturmian words with slope $\alpha$ and we fully characterise their balance properties in terms of the Ostrowski representations of $m$ an $n$ with respect to $\alpha$. This generalises recent results by Anselmo et al., as well as by Shallit and Vukusic, where only quadratic irrationals $\alpha$ were considered. In contrast to the two mentioned papers, our approach is based on the distribution of $n\alpha\bmod 1$.
@@ 1.  Introduction
For example, setting $\alpha=(\sqrt{5}-1)/2$ and $\theta=0$ we get the famous infinite Fibonacci word
$\mathbf{f}=101101011011010\cdots.$
(Note that in the common definition of the Fibonacci word the 0’s and $1$’s are flipped, but we stick with our definition for consistency.)
Recall that a factor is simply a contiguous block of symbols within a word, and the weight of a factor of a binary word is the number of contained $1$’s. One of the basic properties of Sturmian words is that they are balanced, that is, the weight of any two factors of the same length differs by at most $1$. (For example, in the Fibonacci word above, all factors of length $5$ either contain exactly three or exactly four $1$’s.) In fact, the converse holds as well: Every non-periodic balanced word is a Sturmian word.
@@
For every word $\mathbf{a}=a_{1}a_{2}a_{3}\cdots$ let us define the infinite Hankel matrix $A=(a_{k,\ell})_{k\geq 1,\ell\geq 0}$ by $a_{k,\ell}:=a_{k+\ell}$. Then we can look at $m\times n$ submatrices where the entry in the upper left corner has index sum $i$:
@@
Let $\mathbf{a}=a_{1}a_{2}\cdots$ be an infinite word over $\{0,1\}$. We say that the $m\times n$ rectangles of $\mathbf{a}$ are balanced if there exists an integer $c=c(\mathbf{a},m,n)$ such that
@@
for all $i\geq 1$.
For example, for the Fibonacci word it turns out that the $2\times 3$ rectangles are balanced (they always have either weight $3$ or $4$), whereas the $2\times 4$ rectangles are not balanced (they can have weight $4$, 55, or $6$). Anselmo et al. [2] proved that for $m,n\geq 2$, if $\max(m,n)$ is a Fibonacci number, then the $m\times n$ rectangles of the Fibonacci word are balanced. In [7] a full characterisation of balancedness was given in terms of the Zeckendorf representations of $m,n$ (that is, the unique representation of $m,n$ as the sum of distinct non-consecutive Fibonacci numbers). Moreover, it was described how the software Walnut can be used to do the same for every fixed quadratic irrational $\alpha$ and the corresponding representations. Note that assuming $\alpha$ is a quadratic irrational is essential for using Walnut because quadratic irrationals are exactly the numbers with periodic continued fraction expansion.
In this paper, we completely solve the two dimensional balance problem for all irrational $\alpha$. The characterisation (Theorem 2.1 in the next section) is in terms of the Ostrowski representations of $m,n$ with respect to $\alpha$. The proof is based on Diophantine approximation and ideas used by Berthé and Tijdeman [4]. Berthé and Tijdeman also mentioned the connection between balancedness and so-called bounded remainder sets, which is a concept from dynamical systems/discrepancy theory. We provide a little bit of background, as our main result will turn out to be equivalent to a specific statement about the distribution of $n\alpha\bmod 1$. For some quick intuition on this, note that  (1.1) can equivalently be phrased as
@@
It is well known (see, e.g., [5, Chapt. 1]) that for irrational $\alpha$ the sequence $(n\alpha)_{n\geq 0}$ is uniformly distributed modulo $1$. In other words, if we consider the fractional parts $\left\{n\alpha\right\}=n\alpha-\lfloor n\alpha\rfloor\in[0,1)$ for $n=0,1,2,\ldots$, every interval $I\subseteq[0,1)$ gets “its fair share of points” in the following sense:
@@
Let $\alpha,d\in(0,1)$ and let $N\geq 1$ be an integer. We say the intervals of length $d$ are balanced with respect to $(\alpha,N)$ if there exists an integer $c=c(\alpha,d,N)$ such that for all half open intervals $I=[x,x+d)$, $0\leq x<1$, we have
@@
Note that the intervals $I$ are understood modulo $1$, i.e., if $x+d\geq 1$ then the interval is $I=[x,1)\cup[0,\left\{x+d\right\})$.
@@
In the next section, we state the full characterisation of balanced $m\times n$ rectangles. In Section 3 we prove the equivalence between balanced rectangles and balanced intervals. Then we make preparations for proving the characterisation of balanced intervals of length $\left\{n\alpha\right\}$: In Section 4 we rephrase the balancedness in terms of the bijectiveness of a certain function; in Section 5 we recall some specific properties of Ostrowski representations. Finally, in Section 6, we prove the full characterisation of balanced intervals, and thus of balanced $m\times n$ rectangles.
We conclude this introduction with a remark on the intercept $\theta$ of Sturmian words.
@@
In the rest of the paper, we will assume $\theta=0$. This can be justified by the following. Recall that the factors of a word are all the finite contiguous blocks of symbols within the word. It is a basic fact (see, e.g., [1, Theorem 10.5.3]) that two Sturmian words have the same slope if and only if they have the same sets of factors. Since each rectangle $A(i,m,n)$ is fully determined by the factor $a_{i}\cdots a_{i+m+n-2}$, the infinite Hankel matrices corresponding to two Sturmian words with the same slope have exactly the same rectangles.
2.  Full characterization via Ostrowski representations
In this section we state our main result, namely the full characterisation of the $m\times n$ balanced rectangles. But first, let us recall continued fractions and the Ostrowski representation; see, e.g., [6] for a reference book.
@@ 2.  Full characterization via Ostrowski representations
where $a_{0}$ is an integer and $a_{1},a_{2},\ldots$ are positive integers, called partial quotients. We can truncate the continued fraction expansion of $\alpha$ at its $k$-th partial quotient, and obtain the rational number $p_{k}/q_{k}=[a_{0};a_{1},\ldots,a_{k}]$, called the $k$-th convergent to $\alpha$. These convergents are famously particularly good approximations to $\alpha$. In particular, the numbers $q_{k}\alpha$ are very close to an integer (at least for large $k$), or, in other words, very close to 0 modulo $1$. Since we are interested in $n\alpha\bmod 1$, the natural way to represent an integer $n$ therefore is to write it as the sum of $q_{k}$’s. This is known as the Ostrowski representation. To be precise, every positive integer $n$ has a unique representation
@@
Now we state our characterisation of balanced $m\times n$ rectangles in terms of the Ostrowski representations of $m,n$. Since the infinite Hankel matrix $A=(a_{i,j})_{i\geq 0,j\geq 0}$ is symmetric, the balance problem for rectangles is symmetric, and from now on we assume $m\leq n$. Moreover, we can assume $m\geq 2$ because for $m=1$ the rectangles are just the factors of the one-dimensional word, which are of course balanced. It turns out that there are essentially only two situations when the rectangles with $2\leq m\leq n$ are balanced:
  • The integer $m$ only has small digits, and $n$ only has large digits in its Ostrowski representation. In the edge case, where $m$ and $n$ share exactly one digit, there are some extra conditions.
@@
Let $\alpha\in(0,1)$ be irrational and $2\leq m\leq n$. Then the $m\times n$ rectangles of the Sturmian words with slope $\alpha$ are balanced if and only if the Ostrowski representations of $m,n$ with respect to $\alpha$ are of at least one of the following four shapes:
@@
For $\alpha=\pi/4$ we have $\alpha=[0;1,3,1,1,1,15,2,72,\ldots]$, and $q_{0}=1,q_{1}=1,q_{2}=4,q_{3}=5,q_{4}=9,\ldots$. Thus, if we want to know whether the $2\times 10$ rectangles are balanced, we can use a greedy algorithm to get the Ostrowski representations
@@ 3.  Correspondence between balancedness and low discrepancy intervals
In this section we prove the equivalence between balancedness of rectangles and balancedness of intervals. We start by checking the well known fact that Sturmian words are balanced, as the corresponding formula will be useful in a moment. We can directly compute the weight of a factor of lenght $n$ starting at index $i$ by using (1.1) with $\theta=0$ and telescoping:
@@
Now we can use this to compute $T(i,m,n)$ (defined in (1.2)) by summing over the rows of $A(i,m,n)$. Note that the conditional expression in (3.1) can be expressed using the indicator function
@@
takes exactly two values for all $i$. This and the fact that $i\alpha\bmod 1$ is dense in $[0,1]$ lead to the following theorem. (Recall Definition 1.2 for the balancedness of intervals.)
@@
For the reversed implication let $x,x+\left\{n\alpha\right\})$ be an arbitrary interval of length $\left\{n\alpha\right\}$. If $x,x+\left\{n\alpha\right\})$ contains none of the points $\left\{\ell\alpha\right\}$ with $0\leq\ell\leq m-1$, set $\varepsilon_{1}:=1$. Otherwise, set
$\varepsilon_{1}:=\min_{0\leq\ell\leq m-1}\left\{x+\left\{n\alpha\right\}-\left\{\ell\alpha\right\}\right\},$
@@
$\varepsilon_{2}:=\min_{\begin{subarray}{c}0\leq\ell\leq m-1\\ \left\{\ell\alpha\right\}\neq x\end{subarray}}\left\{x-\left\{\ell\alpha\right\}\right\},$
i.e., $\varepsilon_{2}$ is the distance between the right endpoints of the interval and the closest point $\left\{\ell\alpha\right\}$ lying strictly left of the interval. Set $\varepsilon:=\min\{\varepsilon_{1},\varepsilon_{2}\}$ and recall that $i\alpha\bmod 1$ is dense in $[0,1]$. Thus, there exists an $i\geq 1$ such that $1-\left\{n\alpha\right\}-\left\{i\alpha\right\}\in(x-\varepsilon,x]$. Then, by construction, the interval $x,x+\left\{n\alpha\right\})$ contains exactly the same points as the interval $1-\left\{n\alpha\right\}-\left\{i\alpha\right\},1-\left\{i\alpha\right\})$. Therefore, the balancedness of the $m\times n$ rectangles implies the balancedness of the intervals of length $\left\{n\alpha\right\}$. ∎
@@
Since the infinite Hankel matrix $A=(a_{k,\ell})_{k\geq 1,\ell\geq 0}$ is symmetric, the balance problem for rectangles is symmetric. By Theorem 3.1 we immediately get the following fact: Let $m,n\geq 1$ be integers. Then the intervals of length $\left\{m\alpha\right\}$ are balanced with respect to $(\alpha,n)$ if and only if the intervals of length $\left\{n\alpha\right\}$ are balanced with respect to $(\alpha,m)$.
@@
In view of the above, some of the cases of Theorem 2.1 become quite obvious, provided one knows the basic properties of Ostrowski representations. For example, the cases (i) and (ii) correspond to the intervals of length $\left\{n\alpha\right\}$ being extremely short (or extremely long, but then one can think of the complements). In fact, it is not very hard to prove that in the cases (i), (ii) each interval contains at most one point, and thus the intervals are balanced.
For other cases, one can use other known tricks as well. For example, if $m=q_{M}$, the points $\left\{\ell\alpha\right\}$, for $0\leq\ell\leq m-1$, are particularly evenly distributed and one can use the trick that $\alpha\approx p_{M}/q_{M}$, and therefore $\left\{\ell\alpha\right\}\approx(\ell p_{M}\bmod q_{M})/q_{M}$, to characterise balancedness.
Overall, one can say that the intervals are balanced in two cases: Either the intervals are very short (or very long); this corresponds to the cases (i) and (ii) in Theorem 2.1. Or the points are very evenly distributed, and the interval lengths are either slightly shorter or slightly longer than the distance between certain two points; this corresponds to the cases (iii), (iv) in Theorem 2.1.
In this paper, we want to deal with all cases in a somewhat uniform way: We rephrase balancedness of intervals in terms of bijectivity of a certain map $f$. This is done in the next section. Later, we rephrase the bijectivity of $f$ again in terms of another function, which is more closely related to approximation properties of Ostrowski representations.

### v2

@@ Abstract.
We consider $m\times n$ rectangular matrices formed from Sturmian words with slope $\alpha$, and we fully characterise their balance properties in terms of the Ostrowski representations of $m$ and $n$ with respect to $\alpha$. This generalises recent results by Anselmo et al., as well as those by Shallit and the author, where only quadratic irrational slopes were considered. In contrast to the two mentioned papers, the approach in this paper is based on the distribution of $n\alpha\bmod 1$.
@@ 1.  Introduction
For example, setting $\alpha=(3-\sqrt{5})/2$ and $\theta=0$ we get the famous infinite Fibonacci word
$\mathbf{f}=010010100100101\cdots.$
Recall that a factor is simply a contiguous block of symbols within a word, and the weight of a factor of a binary word is the number of $1$’s contained in it. One of the basic properties of Sturmian words is that they are balanced, that is, the weights of any two factors of the same length differ by at most $1$. (For example, in the Fibonacci word above, all factors of length $5$ either contain exactly one or exactly two $1$’s.) In fact, the converse holds as well: every non-periodic balanced word is a Sturmian word.
@@
For every infinite word $\mathbf{a}=a_{1}a_{2}a_{3}\cdots$ let us define the infinite Hankel matrix $A=(a_{k,\ell})_{k\geq 1,\ell\geq 0}$ by $a_{k,\ell}:=a_{k+\ell}$. Then we can consider $m\times n$ submatrices where the entry in the upper left corner has index sum $i$:
@@
Let $\mathbf{a}=a_{1}a_{2}a_{3}\cdots$ be an infinite word over $\{0,1\}$. We say that the $m\times n$ rectangles of $\mathbf{a}$ are balanced if there exists an integer $c=c(\mathbf{a},m,n)$ such that
@@
for all $i\geq 1$. [^1] This notion of balance for rectangles appears in [6], described via “abelian complexity”.
For example, for the Fibonacci word it turns out that the $2\times 3$ rectangles are balanced (they always have weight $2$ or $3$), whereas the $2\times 4$ rectangles are not balanced (they can have weight $2$, 33, or $4$). Anselmo et al. [2] proved that if $\max(m,n)$ is a Fibonacci number, then the $m\times n$ rectangles of the Fibonacci word are balanced. In [8], a full characterisation of balancedness was given in terms of the Zeckendorf representations of $m,n$. (The Zeckendorf representation of a positive integer is its unique representation as the sum of distinct and non-consecutive Fibonacci numbers). Moreover, it was described how the software Walnut can be used to do the same for every fixed quadratic irrational $\alpha$ and the corresponding representations. Note that the assumption that $\alpha$ is a quadratic irrational is essential for using Walnut, since quadratic irrationals are precisely the numbers with eventually periodic continued fraction expansion.
In this paper, we completely solve the $2$-dimensional balance problem for all irrationals $\alpha$. The characterisation (Theorem 2.1 in the next section) is in terms of the Ostrowski representations of $m,n$ with respect to $\alpha$. The proof is based on Diophantine approximation and ideas used by Berthé and Tijdeman [4]. Berthé and Tijdeman also mentioned the connection between balance and so-called bounded remainder sets, which is a concept from dynamical systems/discrepancy theory. We provide a little bit of background, as our main result will turn out to be equivalent to a specific statement about the distribution of $n\alpha\bmod 1$. For some quick intuition on this, note that  (1.1) can equivalently be phrased as
@@
It is well known (see, e.g., [5, Chapt. 1]) that for irrational $\alpha$ the sequence $(n\alpha)_{n\geq 0}$ is uniformly distributed modulo $1$. In other words, if we consider the fractional parts $\left\{n\alpha\right\}$ for $n=0,1,2,\ldots$, every interval $I\subseteq[0,1)$ gets “its fair share of points” in the following sense:
@@
Let $\alpha,\delta\in(0,1)$ and let $N\geq 1$ be an integer. We say the intervals of length $\delta$ are balanced with respect to $(\alpha,N)$ if there exists an integer $c=c(\alpha,\delta,N)$ such that for all half open intervals $I=[\xi,\xi+\delta)$, $0\leq\xi<1$, we have
@@
Note that the intervals $I$ are understood modulo $1$, i.e., if $\xi+\delta\geq 1$ then the interval is $I=[\xi,1)\cup[0,\left\{\xi+\delta\right\})$.
@@
In the next section, we state the full characterisation of balanced $m\times n$ rectangles. In Section 3 we prove the equivalence between balanced rectangles and balanced intervals. Then we make preparations for proving the characterisation of balanced intervals of length $\left\{n\alpha\right\}$: In Section 4 we rephrase balancedness in terms of bijectivity of a certain function; in Section 5 we recall some specific properties of Ostrowski representations. Finally, in Section 6, we prove the full characterisation of balanced $m\times n$ rectangles.
We conclude this introduction with two remarks on the parameters $\alpha$ and $\theta$ of Sturmian words.
@@
In the rest of the paper, we will assume $\theta=0$. This can be justified by the following: It is a basic fact (see, e.g., [1, Theorem 10.5.3]) that two Sturmian words have the same slope if and only if they have the same sets of factors. Since each rectangle $A(i,m,n)$ is fully determined by the factor $a_{i}\cdots a_{i+m+n-2}$, the infinite Hankel matrices corresponding to two Sturmian words with the same slope have exactly the same rectangles.
Remark 1.4.
We will also often assume $\alpha<1/2$. It is easy to check that the Sturmian word with slope $1-\alpha$ and intercept 0 can be obtained by flipping the digits of the Sturmian word with slope $\alpha$ and intercept 0. Of course, for every pair $(m,n)$, the $m\times n$ rectangles of one sequence are balanced if and only if the $m\times n$ rectangles of the other sequence are. Therefore, if $\alpha>1/2$, we can equivalently consider $1-\alpha<1/2$ instead.
2.  Full characterisation via Ostrowski representations
In this section we state our main result, namely the full characterisation of the $m\times n$ balanced rectangles. But first, let us briefly recall continued fractions and the Ostrowski representation; see, e.g., [7] for a reference book.
@@ 2.  Full characterisation via Ostrowski representations
where $a_{0}$ is an integer and $a_{1},a_{2},\ldots$ are positive integers, called partial quotients. We can truncate the continued fraction expansion of $\alpha$ at its $k$-th partial quotient, and obtain the rational number $p_{k}/q_{k}=[a_{0};a_{1},\ldots,a_{k}]$, called the $k$-th convergent to $\alpha$. These convergents are famously particularly good approximations to $\alpha$. In particular, the numbers $q_{k}\alpha$ are very close to an integer (at least for large $k$), or, in other words, very close to 0 modulo $1$. Since we are interested in $n\alpha\bmod 1$, the natural way to represent an integer $n$ therefore is to write it as the sum of $q_{k}$’s, using a greedy algorithm. This is known as the Ostrowski representation. To be precise, every positive integer $n$ has a unique representation
@@
Now we state our characterisation of balanced $m\times n$ rectangles of a Sturmian word with slope $\alpha$ in terms of the Ostrowski representations of $m,n$. Since the infinite Hankel matrix $A=(a_{k,\ell})_{k\geq 1,\ell\geq 0}$ is symmetric, the balance problem for rectangles is symmetric, and from now on we assume $m\leq n$. Moreover, we can assume $m\geq 2$ because for $m=1$ the rectangles are just the factors of the $1$-dimensional Sturmian word, which are of course balanced. It turns out that there are essentially only two situations when the rectangles with $2\leq m\leq n$ are balanced:
  • The integer $m$ only has small digits, and $n$ only has large digits in its Ostrowski representation with respect to $\alpha$. In the edge case, where $m$ and $n$ share exactly one digit, there are some extra conditions.
@@
Let $\alpha\in(0,1)$ be irrational and $2\leq m\leq n$. Then the $m\times n$ rectangles of the Sturmian words with slope $\alpha$ are balanced if and only if the Ostrowski representations of $m,n$ with respect to $\alpha$ are of at least one of the following four shapes.
@@
For $\alpha=\pi/4$ we have $\alpha=[0;1,3,1,1,1,15,2,72,\ldots]$, and $q_{0}=1$, $q_{1}=1$, $q_{2}=4$, $q_{3}=5$, $q_{4}=9$, …. Thus, if we want to know whether the $2\times 10$ rectangles are balanced, we can use a greedy algorithm to get the Ostrowski representations
@@
Remark 2.3.
In Remark 1.4 we mentioned that from a balance point of view it doesn’t matter if we consider $\alpha$ or $1-\alpha$. Indeed, the same is true for the shape of the Ostrowski representations. The next lemma implies, in particular, that the Ostrowski representations of $m,n$ with respect to $\alpha$ are of one of the four special shapes in Theorem 2.1 if and only if those with respect to $1-\alpha$ are. Together with Remark 1.4, this allows us to assume $\alpha<1/2$ in the rest of the paper, without loss of generality.
Note that for $\alpha=[0;a_{1},a_{2},\ldots]$ we have $\alpha>1/2\iff a_{1}=1$, in which case, by the rules of the Ostrowski representation, we have $b_{0}=0$. In other words, if $\alpha>1/2$, then the first possible nonzero term in the representation is $b_{1}q_{1}$.
Lemma 2.4.
Let $\alpha<1/2$ be irrational and $n$ a positive integer. Then the Ostrowski representation of $n$ with respect to $\alpha$ is $n=\sum_{k=0}^{N}b_{k}q_{k}$ if and only if the Ostrowski representation of $n$ with respect to $1-\alpha$ is $n=\sum_{k=1}^{N+1}b_{k-1}q_{k}$.
Proof.
For $\alpha<1/2$ and $\alpha=[0;a_{1},a_{2},a_{3},\ldots]$ one can check that $1-\alpha=[0;1,a_{1}-1,a_{2},a_{3},\ldots]$. Then from the recurrence formula for convergents (see also (5.1) in Section 5.1) one can see that $\alpha$ and $1-\alpha$ have the same sequence of denominators $q_{0},q_{1},q_{2},\ldots$, except for the index shift. ∎
@@ 3.  Correspondence between balancedness and low discrepancy intervals
In this section we prove the equivalence between balancedness of rectangles and balancedness of intervals. We start by checking the well known fact that Sturmian words are balanced, as the corresponding formula will be useful in a moment. We can directly compute the weight of a factor of length $n$ starting at index $i$ by using (1.1) with $\theta=0$ and telescoping:
@@
We can now use this to compute $T(i,m,n)$ (defined in (1.2)) by summing over the rows of $A(i,m,n)$. Note that the conditional expression in (3.1) can be expressed using the indicator function in the following way:
@@
takes exactly two values for all $i$. This and the fact that $i\alpha\bmod 1$ is dense in $[0,1)$ lead to the following theorem. (Recall Definition 1.2 for the balancedness of intervals.)
@@
For the reversed implication let $\xi,\xi+\left\{n\alpha\right\})$ be an arbitrary interval of length $\left\{n\alpha\right\}$. If $\xi,\xi+\left\{n\alpha\right\})$ contains none of the points $\left\{\ell\alpha\right\}$ with $0\leq\ell\leq m-1$, set $\varepsilon_{1}:=1$. Otherwise, set
$\varepsilon_{1}:=\min_{0\leq\ell\leq m-1}\left\{\xi+\left\{n\alpha\right\}-\left\{\ell\alpha\right\}\right\},$
@@
$\varepsilon_{2}:=\min_{\begin{subarray}{c}0\leq\ell\leq m-1\\ \left\{\ell\alpha\right\}\neq\xi\end{subarray}}\left\{\xi-\left\{\ell\alpha\right\}\right\},$
i.e., $\varepsilon_{2}$ is the distance between the left endpoint of the interval and the closest point $\left\{\ell\alpha\right\}$ lying strictly to the left of the interval. Set $\varepsilon:=\min\{\varepsilon_{1},\varepsilon_{2}\}$ and recall that $i\alpha\bmod 1$ is dense in $[0,1)$. Thus, there exists an $i\geq 1$ such that $1-\left\{n\alpha\right\}-\left\{i\alpha\right\}\in(\xi-\varepsilon,\xi]$. Then, by construction, the interval $\xi,\xi+\left\{n\alpha\right\})$ contains exactly the same points as the interval $1-\left\{n\alpha\right\}-\left\{i\alpha\right\},1-\left\{i\alpha\right\})$. Therefore, the balancedness of the $m\times n$ rectangles implies the balancedness of the intervals of length $\left\{n\alpha\right\}$. ∎
@@
As mentioned before, since the infinite Hankel matrix $A=(a_{k,\ell})_{k\geq 1,\ell\geq 0}$ is symmetric, the balance problem for rectangles is symmetric. From Theorem 3.1 we immediately get the following fact: Let $m,n\geq 1$ be integers. Then the intervals of length $\left\{m\alpha\right\}$ are balanced with respect to $(\alpha,n)$ if and only if the intervals of length $\left\{n\alpha\right\}$ are balanced with respect to $(\alpha,m)$.
@@
In view of Theorem 3.1, some of the cases of Theorem 2.1 become quite obvious, provided one is familiar with the basic properties of Ostrowski representations. For example, the cases (i) and (ii), where $n$ has only large digits, correspond to the intervals of length $\left\{n\alpha\right\}$ being extremely short (or extremely long, but then one can think of the complements). In fact, it is not very hard to prove that in the cases (i) and (ii) the corresponding intervals each contain at most one point, and thus they are balanced.
For some of the other cases, one can use other known tricks as well. For example, if $m=q_{M}$, the points $\left\{\ell\alpha\right\}$, for $0\leq\ell\leq m-1$, are particularly evenly distributed and one can use the trick that $\alpha\approx p_{M}/q_{M}$, and therefore $\left\{\ell\alpha\right\}\approx(\ell p_{M}\bmod q_{M})/q_{M}$, to characterise balancedness.
Overall, one can say that the intervals are balanced in the two following cases: Either the intervals are very short (or very long); this corresponds to the cases (i) and (ii) in Theorem 2.1. Or the points are very evenly distributed, and the interval lengths are slightly longer (or slightly shorter) than the distance between certain two points; this corresponds to the cases (iii) and (iv) in Theorem 2.1.
In this paper we want to deal with all cases in a somewhat uniform way. To that aim, we rephrase balancedness of intervals in terms of bijectivity of a certain map $f_{*}$. This is done in the next section. Later, we will rephrase the bijectivity of $f_{*}$ again in terms of another function, which is more closely related to approximation properties of Ostrowski representations.